OpenStudy (anonymous):

an estimate for the square root of -72 is a -8.1 b -8.5 c -9.1 d -9.5

OpenStudy (anonymous):

an estimate for -72 is ____?where is the rest of the question? a -8.1 b -8.5 c -9.1 d -9.5

OpenStudy (anonymous):

well, if I had to guess I would say -8.1, becuase 72/9 = 8...

OpenStudy (anonymous):

A negative number does not have a square root! Unless you consider imaginary numbers..which none of your answer choices are.

OpenStudy (jdoe0001):

\(\large -\sqrt{72}\qquad or\qquad \sqrt{-72}\quad ?\)

OpenStudy (smokeysthename):

if it were to be \[-\sqrt{72}\] than it would be approx. 8.5

OpenStudy (anonymous):

thank you smokeysthename

OpenStudy (anonymous):

Use the fact that 72 is between 8^2 and 9^2

OpenStudy (smokeysthename):

@DemolisionWolf they changed the question @jared101 no prob

OpenStudy (anonymous):

ya.. i'm going to balete my comment now...

OpenStudy (smokeysthename):

btw if it were to be\[\sqrt{-72}\] then it would be approx. 8.5i

OpenStudy (anonymous):

but what if the negative sign was not in the √

OpenStudy (smokeysthename):

then you would apply the negative sign after you found the square root of 72 so in the end the number you found would be negative

OpenStudy (smokeysthename):

which in this case is 8.5 so it would be -8.5